ethat a market is described by the and demand equations: lowng supply and demaned a QP = 300-P e for the equilibrium price and the equilibrium quantity Spose that a tax of T is placed on buyers so the new demand equation is QD = 300-(P + T)
Answer #1

(a) In equilibrium, Qd = Qs.

300 - P = 2P

3P = 300

P = 100

P = 2 x 100 = 200

(b) After tax, equating new demand function with original supply function,

300 - P - T = 2P

3P = 300 - T

P = (300 - T) / 3 [Price received by producers = Market price]

Price paid by buyers = [(300 - T) / 3] + T = (300 - T + 3T) / 3 = (300 + 2T) / 3

Q = 2 x Market price = 2 x [(300 - T) / 3] = (600 - 2T) / 3

Tax burden of buyers = Price paid by buyers after tax - pre-tax price = [(300 + 2T) / 3] - 200 = (300 + 2T - 600) / 3

= (2T - 300) / 3

Tax burden of sellers = Unit tax - Tax burden of buyers = T - [(2T - 300) / 3] = (3T - 2T + 300) / 3 = (T + 300) / 3